OpenStudy (anonymous):

Solve by factoring and using the principle of zero products (5m+7)(m+5)=23

OpenStudy (anonymous):

figured -6-2/5

OpenStudy (swissgirl):

ok so we gotta bring the 23 or over to the other side (5m+7)(m+5)-23=0 Then we foil \(5m^2+25m+7m+35-23=0\) \(5m^2+32m+12=0\)